OpenStudy (anonymous):

A system of equations is shown below: -3x+7y=-16 -9x+5y=16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations. I'm really confused if anyone knows how to do this please help me!!!

OpenStudy (anonymous):

I'm not actually sure about what they're asking us to do here, but if you're going to solve these equations like other simultaneous equations then we can do it like this: Let's denote the equations like so: (1) -3x+7y=-16 (2) -9x+5y=16 Then subtract 3 times (1) to 2 and we get: -9x+5y-3(-3x+7y)=16-3(-16) => -9x+5y+9x-21y=16+48 => -16y=64 => y=-4 Sub y=-4 into (1) or (2): -3x+7y=-16 => -3x+7(-4)=-16 => -3x-28=-16 => -3x=12 => x=-4 And there we have it: x=-4, y=-4. For B I think you just need to put these values of x and y back into (1) and (2).

OpenStudy (anonymous):

Oh wow that makes so much sense now thank you so much!!!!!